# Measuring Vision-Language STEM Skills of Neural Models

### v1

@@ Appendix D Summary of Skills
Subject: Math Skill: absolute-value-and-integers-word-problems Description: Debbie likes watching the show Engineering Marvels. In last night’s episode, the engineering team visited a tall skyscraper and a deep mine. A banner at the bottom of the screen showed the elevation of each location the team visited.Which location is closer to sea level? Picture: [Uncaptioned image] Choices: [bottom of the mine, top of the skyscraper, ] Answer index: 0
Subject: Math Skill: acute-obtuse-and-right-triangles Description: What kind of triangle is this? Picture: [Uncaptioned image] Choices: [obtuse, right, acute, ] Answer index: 0
Subject: Math Skill: acute-right-obtuse-and-straight-angles Description: Is this angle acute, right, obtuse, or straight? Picture: [Uncaptioned image] Choices: [straight, obtuse, acute, right, ] Answer index: 1
Subject: Math Skill: addition-sentences-up-to-10-what-does-the-model-show Description: Which addition sentence does the picture show? Picture: [Uncaptioned image] Choices: [4+3=7, 5+2=7, ] Answer index: 0
Subject: Math Skill: addition-sentences-up-to-10-which-model-matches Description: Which shows 8+1=9? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: addition-sentences-up-to-5-what-does-the-model-show Description: Which addition sentence does the picture show? Picture: [Uncaptioned image] Choices: [4+1=5, 3+1=4, ] Answer index: 0
Subject: Math Skill: addition-sentences-up-to-5-which-model-matches Description: Which shows 2+1=3? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: addition-sentences-using-number-lines-sums-up-to-20 Description: Which addition sentence does this model show? Picture: [Uncaptioned image] Choices: [3+3=6, 5+6=11, 5+4=9, 5+3=8, ] Answer index: 3
@@
Subject: Math Skill: adjust-a-budget Description: This table shows Angie’s February budget.What could Angie do to balance her budget? Picture: [Uncaptioned image] Choices: [spend only \$60 at the mall, and teach a baking class for \$30, decorate more custom cookies to earn another \$35, and spend only \$60 at the mall, spend \$20 more on the baking kit, and teach a baking class for \$30, decorate more custom cookies to earn another \$35, and spend \$20 more on the baking kit, ] Answer index: 1
@@
Subject: Math Skill: am-or-pm Description: Farmer Keenan is getting up to go milk his cows. It is just before sunrise. His watch shows:What time is it? Picture: [Uncaptioned image] Choices: [4:30 P.M., 4:30 A.M., ] Answer index: 1
Subject: Math Skill: angle-angle-criterion-for-similar-triangles Description: FGH and JKL are shown below.Which statement is true? Picture: [Uncaptioned image] Choices: [FGH is similar to JKL., FGH is not similar to JKL., There is not enough information to determine whether the triangles are similar., ] Answer index: 2
Subject: Math Skill: angles-as-fractions-of-a-circle Description: What fraction of the circle does this angle cut out? Picture: [Uncaptioned image] Choices: [1/4, 3/4, 1 whole, 1/2, ] Answer index: 0
@@
Subject: Math Skill: count-money-pennies-and-nickels Description: How much money is there? Picture: [Uncaptioned image] Choices: [6¢, 11¢, 16¢, ] Answer index: 1
Subject: Math Skill: count-money-pennies-through-dimes Description: How much money is there? Picture: [Uncaptioned image] Choices: [18¢, 16¢, 19¢, ] Answer index: 0
Subject: Math Skill: count-on-ten-frames-up-to-10 Description: How many dots are on the frame? Picture: [Uncaptioned image] Choices: [4, 9, 2, 8, 1, 7, 5, 10, 6, 3, ] Answer index: 1
Subject: Math Skill: count-on-ten-frames-up-to-3 Description: How many dots are on the frame? Picture: [Uncaptioned image] Choices: [3, 2, 1, ] Answer index: 1
Subject: Math Skill: count-on-ten-frames-up-to-5 Description: How many squares are on the frame? Picture: [Uncaptioned image] Choices: [4, 5, 1, 2, 3, ] Answer index: 1
Subject: Math Skill: count-pennies Description: How much money is there? Picture: [Uncaptioned image] Choices: [7¢, 6¢, , ] Answer index: 2
Subject: Math Skill: count-pictures-up-to-10 Description: How many parrots are there? Picture: [Uncaptioned image] Choices: [1, 4, 3, 6, 9, 5, 7, 10, 8, 2, ] Answer index: 4
Subject: Math Skill: count-pictures-up-to-3 Description: How many butterflies are there? Picture: [Uncaptioned image] Choices: [2, 1, 3, ] Answer index: 2
Subject: Math Skill: count-pictures-up-to-5 Description: How many snowmen are there? Picture: [Uncaptioned image] Choices: [2, 5, 1, 4, 3, ] Answer index: 4
Subject: Math Skill: count-scattered-shapes-up-to-10 Description: How many shapes are there? Picture: [Uncaptioned image] Choices: [2, 3, 8, 1, 9, 10, 4, 6, 7, 5, ] Answer index: 5
Subject: Math Skill: count-scattered-shapes-up-to-5 Description: How many rectangles are there? Picture: [Uncaptioned image] Choices: [2, 4, 5, 3, 1, ] Answer index: 2
Subject: Math Skill: count-shapes-in-rings-up-to-10 Description: How many triangles are there? Picture: [Uncaptioned image] Choices: [9, 3, 5, 2, 7, 6, 8, 1, 4, 10, ] Answer index: 4
Subject: Math Skill: count-shapes-in-rows-up-to-10 Description: How many hearts are there? Picture: [Uncaptioned image] Choices: [5, 1, 7, 9, 8, 10, 3, 4, 2, 6, ] Answer index: 3
Subject: Math Skill: count-shapes-in-rows-up-to-5 Description: How many shapes are there? Picture: [Uncaptioned image] Choices: [3, 2, 1, 4, 5, ] Answer index: 0
Subject: Math Skill: count-shapes-up-to-3 Description: How many triangles are there? Picture: [Uncaptioned image] Choices: [2, 1, 3, ] Answer index: 0
Subject: Math Skill: count-sides Description: Which shape has 4 sides? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Math Skill: count-sides-and-corners Description: Which shape has 5 corners? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Math Skill: count-to-100 Description: How many dots are there? Picture: [Uncaptioned image] Choices: [46, 49, 44, ] Answer index: 0
@@
Subject: Math Skill: identify-alternate-interior-and-alternate-exterior-angles Description: line{RT} and line{UW} are parallel lines.Which angles are alternate interior angles? Picture: [Uncaptioned image] Choices: [angle{TSV} and angle{UVS}, angle{TSV} and angle{TSQ}, angle{TSV} and angle{RSV}, angle{TSV} and angle{WVS}, ] Answer index: 0
@@
Subject: Math Skill: identify-complementary-supplementary-vertical-adjacent-and-congruent-angles Description: Which angle is vertical to angle{3}? Picture: [Uncaptioned image] Choices: [angle{6}, angle{5}, angle{4}, angle{2}, ] Answer index: 0
Subject: Math Skill: identify-complementary-supplementary-vertical-and-adjacent-angles Description: Which angles are adjacent to each other? Picture: [Uncaptioned image] Choices: [angle{1}angle{3} and angle{7}, angle{1}angle{5} and angle{1}angle{4}, angle{8} and angle{4}, angle{1}angle{0} and angle{4}, ] Answer index: 1
Subject: Math Skill: identify-congruent-figures Description: Are these shapes congruent? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 1
Subject: Math Skill: identify-direct-variation-and-inverse-variation Description: Which equation shows direct variation? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Math Skill: identify-equivalent-expressions-using-strip-models Description: This model represents the expression x+x+1+1.Which expression is equivalent to x+x+1+1? Picture: [Uncaptioned image] Choices: [4x, 2x+3, 3x+1, 2x+2, ] Answer index: 3
Subject: Math Skill: identify-equivalent-fractions-on-number-lines Description: Is 1/2 equivalent to 1/3 ? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 0
Subject: Math Skill: identify-equivalent-fractions-using-number-lines Description: Is 2/3 equivalent to 4/6 ? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 0
Subject: Math Skill: identify-equivalent-linear-expressions-using-algebra-tiles Description: These tiles represent the expression 3x+5x.Which expression is equivalent to 3x+5x? Picture: [Uncaptioned image] Choices: [x+8, 2x, 8x, 8x+2, ] Answer index: 2
Subject: Math Skill: identify-faces-of-three-dimensional-figures Description: Which shape has a circle as a face? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Math Skill: identify-faces-of-three-dimensional-shapes Description: Which shape has a circle as a face? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Math Skill: identify-fourths Description: Look at the colored part of each shape. Which shape shows one-fourth? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
Subject: Math Skill: identify-functions Description: Look at this graph:Is this relation a function? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 1
Subject: Math Skill: identify-functions-graphs Description: Which of these relations is a function? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
Subject: Math Skill: identify-functions-vertical-line-test Description: Which of these relations is a function? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 3
Subject: Math Skill: identify-graphs-of-continuous-functions Description: Is the function f(x) continuous? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 0
@@
Subject: Math Skill: identify-halves Description: Look at the colored part of each shape. Which shape shows one-half? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: identify-halves-and-fourths Description: Which figure shows fourths? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 1
@@
Subject: Math Skill: lines-line-segments-and-rays Description: What is this? Picture: [Uncaptioned image] Choices: [line, line segment, ray, ] Answer index: 1
@@
Subject: Math Skill: make-predictions-with-scatter-plots Description: Based on the scatter plot below, which is a better prediction for x when y = 46? Picture: [Uncaptioned image] Choices: [50, 98, ] Answer index: 0
Subject: Math Skill: match-addition-sentences-and-models-sums-to-10 Description: Which shows 2+2=4? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Math Skill: match-analog-and-digital-clocks Description: Look at the analog clock:Which digital clock shows the same time? Picture: [Uncaptioned image] Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: match-analog-clocks-and-times Description: What time does the clock show? Picture: [Uncaptioned image] Choices: [5:00, 4:30, ] Answer index: 0
Subject: Math Skill: match-clocks-and-times Description: What time does the clock show? Picture: [Uncaptioned image] Choices: [eight fifty, seven fifty, nine forty, ] Answer index: 1
Subject: Math Skill: match-digital-clocks-and-times Description: Which clock shows six thirty-five? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 1
Subject: Math Skill: match-exponential-functions-and-graphs Description: formula_desc.png Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: match-exponential-functions-and-graphs-ii Description: formula_desc.png Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Skill: match-fractions-to-models-halves-thirds-and-fourths Description: Look at the colored part of each shape. Which shape shows one-fourth? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 2
Subject: Math Skill: match-mixed-numbers-to-models Description: Which mixed number is shown? Picture: [Uncaptioned image] Choices: [3 3/8, 4 2/8, 3 2/8, 3 5/8, ] Answer index: 0
@@
Subject: Math Skill: mean-find-the-missing-number Description: Susan has the following data:If the mean is 25, which number could r be? Picture: [Uncaptioned image] Choices: [29, 38, ] Answer index: 0
Subject: Math Skill: mean-median-mode-and-range-find-the-missing-number Description: Jayla has the following data:If the mean is 14, which number could s be? Picture: [Uncaptioned image] Choices: [11, 3, ] Answer index: 0
Subject: Math Skill: measure-angles-with-a-protractor Description: Is this angle acute, right, or obtuse? Picture: [Uncaptioned image] Choices: [right, obtuse, acute, ] Answer index: 2
@@
Subject: Math Skill: median-find-the-missing-number Description: Danny has the following data:If the median is 97, which number could c be? Picture: [Uncaptioned image] Choices: [98, 47, ] Answer index: 0
Subject: Math Skill: model-and-solve-equations-using-algebra-tiles Description: Which equation does this set of algebra tiles represent? Picture: [Uncaptioned image] Choices: [– 4x–1= – 9, – 8x–1= – 9, 8x–1= – 9, – x–1= – 10, ] Answer index: 3
Subject: Math Skill: model-and-solve-linear-equations-using-algebra-tiles Description: Which equation does this set of algebra tiles represent? Picture: [Uncaptioned image] Choices: [3x=27, 3x=24, 2x=26, 2x=24, ] Answer index: 1
Subject: Math Skill: more Description: Which group has more? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
@@
Subject: Math Skill: reflection-rotation-and-translation Description: How has this figure been transformed?It has been… Picture: [Uncaptioned image] Choices: [translated, reflected, rotated, ] Answer index: 1
Subject: Math Skill: regular-and-irregular-polygons Description: Is this shape a regular polygon? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 1
Subject: Math Skill: represent-numbers-up-to-10 Description: Which group has 6 triangles? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: represent-numbers-up-to-20 Description: Which picture shows 8 dots? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: represent-numbers-with-pictures-up-to-3 Description: Which shows 2? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: represent-numbers-with-pictures-up-to-5 Description: Which shows 1? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: represent-numbers-with-shapes-up-to-3 Description: Which group has 3 circles? Picture: None Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: represent-numbers-with-shapes-up-to-5 Description: Which group has 4 hexagons? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: rhombuses Description: Which shape is a rhombus? Picture: None Choices: [Uncaptioned image][Uncaptioned image] Answer index: 0
Subject: Math Skill: rotational-symmetry Description: Does this picture have rotational symmetry? Picture: [Uncaptioned image] Choices: [no, yes, ] Answer index: 0
@@
Subject: Math Skill: rotational-symmetry-amount-of-rotation Description: This image has rotational symmetry. What is the smallest fraction of a full turn you need to rotate the image for it to look the same? Picture: [Uncaptioned image] Choices: [1 2 of a full turn, 1 6 of a full turn, 1 4 of a full turn, 1 3 of a full turn, ] Answer index: 0
Subject: Math Skill: scalene-isosceles-and-equilateral-triangles Description: Is this triangle scalene? Picture: [Uncaptioned image] Choices: [yes, no, ] Answer index: 1

### v2

@@ Appendix D Summary of Skills
Description: Debbie likes watching the show Engineering Marvels. In last night’s episode, the engineering team visited a tall skyscraper and a deep mine. A banner at the bottom of the screen showed the elevation of each location the team visited.Which location is closer to sea level?
Picture: [Uncaptioned image]
Choices: [bottom of the mine, top of the skyscraper, ]
Answer index: 0
Subject: Math
Skill: acute-obtuse-and-right-triangles
Description: What kind of triangle is this?
Picture: [Uncaptioned image]
Choices: [obtuse, right, acute, ]
Answer index: 0
Subject: Math
Skill: acute-right-obtuse-and-straight-angles
Description: Is this angle acute, right, obtuse, or straight?
Picture: [Uncaptioned image]
Choices: [straight, obtuse, acute, right, ]
Answer index: 1
Subject: Math
Skill: addition-sentences-up-to-10-what-does-the-model-show
Description: Which addition sentence does the picture show?
Picture: [Uncaptioned image]
Choices: [4+3=7, 5+2=7, ]
Answer index: 0
Subject: Math
Skill: addition-sentences-up-to-10-which-model-matches
Description: Which shows 8+1=9?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: addition-sentences-up-to-5-what-does-the-model-show
Description: Which addition sentence does the picture show?
Picture: [Uncaptioned image]
Choices: [4+1=5, 3+1=4, ]
Answer index: 0
Subject: Math
Skill: addition-sentences-up-to-5-which-model-matches
Description: Which shows 2+1=3?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: addition-sentences-using-number-lines-sums-up-to-20
Description: Which addition sentence does this model show?
Picture: [Uncaptioned image]
Choices: [3+3=6, 5+6=11, 5+4=9, 5+3=8, ]
Answer index: 3
Subject: Math
Skill: adjust-a-budget
Description: This table shows Angie’s February budget.What could Angie do to balance her budget?
Picture: [Uncaptioned image]
@@
Choices: [spend only \$60 at the mall, and teach a baking class for \$30, decorate more custom cookies to earn another \$35, and spend only \$60 at the mall, spend \$20 more on the baking kit, and teach a baking class for \$30, decorate more custom cookies to earn another \$35, and spend \$20 more on the baking kit, ]
Answer index: 1
Subject: Math
Skill: am-or-pm
@@
Description: Farmer Keenan is getting up to go milk his cows. It is just before sunrise. His watch shows:What time is it?
Picture: [Uncaptioned image]
Choices: [4:30 P.M., 4:30 A.M., ]
Answer index: 1
Subject: Math
Skill: angle-angle-criterion-for-similar-triangles
Description: FGH and JKL are shown below.Which statement is true?
Picture: [Uncaptioned image]
Choices: [FGH is similar to JKL., FGH is not similar to JKL., There is not enough information to determine whether the triangles are similar., ]
Answer index: 2
Subject: Math
Skill: angles-as-fractions-of-a-circle
Description: What fraction of the circle does this angle cut out?
Picture: [Uncaptioned image]
Choices: [1/4, 3/4, 1 whole, 1/2, ]
Answer index: 0
@@
Subject: Math
Skill: count-money-pennies-and-nickels
Description: How much money is there?
Picture: [Uncaptioned image]
Choices: [6¢, 11¢, 16¢, ]
Answer index: 1
Subject: Math
Skill: count-money-pennies-through-dimes
Description: How much money is there?
Picture: [Uncaptioned image]
Choices: [18¢, 16¢, 19¢, ]
Answer index: 0
Subject: Math
Skill: count-on-ten-frames-up-to-10
Description: How many dots are on the frame?
Picture: [Uncaptioned image]
Choices: [4, 9, 2, 8, 1, 7, 5, 10, 6, 3, ]
Answer index: 1
Subject: Math
Skill: count-on-ten-frames-up-to-3
Description: How many dots are on the frame?
Picture: [Uncaptioned image]
Choices: [3, 2, 1, ]
Answer index: 1
Subject: Math
Skill: count-on-ten-frames-up-to-5
Description: How many squares are on the frame?
Picture: [Uncaptioned image]
Choices: [4, 5, 1, 2, 3, ]
Answer index: 1
Subject: Math
Skill: count-pennies
Description: How much money is there?
Picture: [Uncaptioned image]
Choices: [7¢, 6¢, , ]
Answer index: 2
Subject: Math
Skill: count-pictures-up-to-10
Description: How many parrots are there?
Picture: [Uncaptioned image]
Choices: [1, 4, 3, 6, 9, 5, 7, 10, 8, 2, ]
Answer index: 4
Subject: Math
Skill: count-pictures-up-to-3
Description: How many butterflies are there?
Picture: [Uncaptioned image]
Choices: [2, 1, 3, ]
Answer index: 2
Subject: Math
Skill: count-pictures-up-to-5
Description: How many snowmen are there?
Picture: [Uncaptioned image]
Choices: [2, 5, 1, 4, 3, ]
Answer index: 4
Subject: Math
Skill: count-scattered-shapes-up-to-10
Description: How many shapes are there?
Picture: [Uncaptioned image]
Choices: [2, 3, 8, 1, 9, 10, 4, 6, 7, 5, ]
Answer index: 5
Subject: Math
Skill: count-scattered-shapes-up-to-5
Description: How many rectangles are there?
Picture: [Uncaptioned image]
Choices: [2, 4, 5, 3, 1, ]
Answer index: 2
Subject: Math
Skill: count-shapes-in-rings-up-to-10
Description: How many triangles are there?
Picture: [Uncaptioned image]
Choices: [9, 3, 5, 2, 7, 6, 8, 1, 4, 10, ]
Answer index: 4
Subject: Math
Skill: count-shapes-in-rows-up-to-10
Description: How many hearts are there?
Picture: [Uncaptioned image]
Choices: [5, 1, 7, 9, 8, 10, 3, 4, 2, 6, ]
Answer index: 3
Subject: Math
Skill: count-shapes-in-rows-up-to-5
Description: How many shapes are there?
Picture: [Uncaptioned image]
Choices: [3, 2, 1, 4, 5, ]
Answer index: 0
Subject: Math
Skill: count-shapes-up-to-3
Description: How many triangles are there?
Picture: [Uncaptioned image]
Choices: [2, 1, 3, ]
Answer index: 0
Subject: Math
Skill: count-sides
Description: Which shape has 4 sides?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: count-sides-and-corners
Description: Which shape has 5 corners?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: count-to-100
Description: How many dots are there?
Picture: [Uncaptioned image]
Choices: [46, 49, 44, ]
Answer index: 0
@@
Subject: Math
Skill: identify-alternate-interior-and-alternate-exterior-angles
Description:
line{RT} and
line{UW} are parallel lines.Which angles are alternate interior angles?
Picture: [Uncaptioned image]
Choices: [angle{TSV} and angle{UVS}, angle{TSV} and angle{TSQ}, angle{TSV} and angle{RSV}, angle{TSV} and angle{WVS}, ]
Answer index: 0
Subject: Math
@@
Skill: identify-complementary-supplementary-vertical-adjacent-and-congruent-angles
Description: Which angle is vertical to angle{3}?
Picture: [Uncaptioned image]
Choices: [angle{6}, angle{5}, angle{4}, angle{2}, ]
Answer index: 0
Subject: Math
Skill: identify-complementary-supplementary-vertical-and-adjacent-angles
Description: Which angles are adjacent to each other?
Picture: [Uncaptioned image]
Choices: [angle{1}angle{3} and angle{7}, angle{1}angle{5} and angle{1}angle{4}, angle{8} and angle{4}, angle{1}angle{0} and angle{4}, ]
Answer index: 1
Subject: Math
Skill: identify-congruent-figures
Description: Are these shapes congruent?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 1
Subject: Math
Skill: identify-direct-variation-and-inverse-variation
Description: Which equation shows direct variation?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: identify-equivalent-expressions-using-strip-models
Description: This model represents the expression x+x+1+1.Which expression is equivalent to x+x+1+1?
Picture: [Uncaptioned image]
Choices: [4x, 2x+3, 3x+1, 2x+2, ]
Answer index: 3
Subject: Math
Skill: identify-equivalent-fractions-on-number-lines
Description: Is 1/2 equivalent to 1/3 ?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 0
Subject: Math
Skill: identify-equivalent-fractions-using-number-lines
Description: Is 2/3 equivalent to 4/6 ?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 0
Subject: Math
Skill: identify-equivalent-linear-expressions-using-algebra-tiles
Description: These tiles represent the expression 3x+5x.Which expression is equivalent to 3x+5x?
Picture: [Uncaptioned image]
Choices: [x+8, 2x, 8x, 8x+2, ]
Answer index: 2
Subject: Math
Skill: identify-faces-of-three-dimensional-figures
Description: Which shape has a circle as a face?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: identify-faces-of-three-dimensional-shapes
Description: Which shape has a circle as a face?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: identify-fourths
@@
Description: Look at the colored part of each shape. Which shape shows one-fourth?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Math
Skill: identify-functions
Description: Look at this graph:Is this relation a function?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 1
Subject: Math
Skill: identify-functions-graphs
Description: Which of these relations is a function?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Math
Skill: identify-functions-vertical-line-test
Description: Which of these relations is a function?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 3
Subject: Math
Skill: identify-graphs-of-continuous-functions
Description: Is the function f(x) continuous?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 0
Subject: Math
Skill: identify-halves
@@
Description: Look at the colored part of each shape. Which shape shows one-half?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: identify-halves-and-fourths
Description: Which figure shows fourths?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 1
@@
Subject: Math
Skill: lines-line-segments-and-rays
Description: What is this?
Picture: [Uncaptioned image]
Choices: [line, line segment, ray, ]
Answer index: 1
Subject: Math
Skill: make-predictions-with-scatter-plots
@@
Description: Based on the scatter plot below, which is a better prediction for x when y = 46?
Picture: [Uncaptioned image]
Choices: [50, 98, ]
Answer index: 0
Subject: Math
Skill: match-addition-sentences-and-models-sums-to-10
Description: Which shows 2+2=4?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: match-analog-and-digital-clocks
@@
Description: Look at the analog clock:Which digital clock shows the same time?
Picture: [Uncaptioned image]
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: match-analog-clocks-and-times
Description: What time does the clock show?
Picture: [Uncaptioned image]
Choices: [5:00, 4:30, ]
Answer index: 0
Subject: Math
Skill: match-clocks-and-times
Description: What time does the clock show?
Picture: [Uncaptioned image]
Choices: [eight fifty, seven fifty, nine forty, ]
Answer index: 1
Subject: Math
Skill: match-digital-clocks-and-times
Description: Which clock shows six thirty-five?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 1
Subject: Math
Skill: match-exponential-functions-and-graphs
Description: formula_desc.png
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: match-exponential-functions-and-graphs-ii
Description: formula_desc.png
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Skill: match-fractions-to-models-halves-thirds-and-fourths
Description: Look at the colored part of each shape. Which shape shows one-fourth?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 2
Subject: Math
Skill: match-mixed-numbers-to-models
Description: Which mixed number is shown?
Picture: [Uncaptioned image]
Choices: [3 3/8, 4 2/8, 3 2/8, 3 5/8, ]
Answer index: 0
Subject: Math
Skill: mean-find-the-missing-number
@@
Description: Susan has the following data:If the mean is 25, which number could r be?
Picture: [Uncaptioned image]
Choices: [29, 38, ]
Answer index: 0
Subject: Math
Skill: mean-median-mode-and-range-find-the-missing-number
Description: Jayla has the following data:If the mean is 14, which number could s be?
Picture: [Uncaptioned image]
Choices: [11, 3, ]
Answer index: 0
Subject: Math
Skill: measure-angles-with-a-protractor
Description: Is this angle acute, right, or obtuse?
Picture: [Uncaptioned image]
Choices: [right, obtuse, acute, ]
Answer index: 2
Subject: Math
Skill: median-find-the-missing-number
@@
Description: Danny has the following data:If the median is 97, which number could c be?
Picture: [Uncaptioned image]
Choices: [98, 47, ]
Answer index: 0
Subject: Math
Skill: model-and-solve-equations-using-algebra-tiles
Description: Which equation does this set of algebra tiles represent?
Picture: [Uncaptioned image]
Choices: [– 4x–1= – 9, – 8x–1= – 9, 8x–1= – 9, – x–1= – 10, ]
Answer index: 3
Subject: Math
Skill: model-and-solve-linear-equations-using-algebra-tiles
Description: Which equation does this set of algebra tiles represent?
Picture: [Uncaptioned image]
Choices: [3x=27, 3x=24, 2x=26, 2x=24, ]
Answer index: 1
Subject: Math
Skill: more
Description: Which group has more?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
@@
Subject: Math
Skill: reflection-rotation-and-translation
Description: How has this figure been transformed?It has been…
Picture: [Uncaptioned image]
Choices: [translated, reflected, rotated, ]
Answer index: 1
Subject: Math
Skill: regular-and-irregular-polygons
Description: Is this shape a regular polygon?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 1
Subject: Math
Skill: represent-numbers-up-to-10
Description: Which group has 6 triangles?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: represent-numbers-up-to-20
Description: Which picture shows 8 dots?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: represent-numbers-with-pictures-up-to-3
Description: Which shows 2?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: represent-numbers-with-pictures-up-to-5
Description: Which shows 1?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: represent-numbers-with-shapes-up-to-3
Description: Which group has 3 circles?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: represent-numbers-with-shapes-up-to-5
Description: Which group has 4 hexagons?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: rhombuses
Description: Which shape is a rhombus?
Picture: None
Choices: [Uncaptioned image][Uncaptioned image]
Answer index: 0
Subject: Math
Skill: rotational-symmetry
Description: Does this picture have rotational symmetry?
Picture: [Uncaptioned image]
Choices: [no, yes, ]
Answer index: 0
Subject: Math
Skill: rotational-symmetry-amount-of-rotation
@@
Description: This image has rotational symmetry. What is the smallest fraction of a full turn you need to rotate the image for it to look the same?
Picture: [Uncaptioned image]
Choices: [1 2 of a full turn, 1 6 of a full turn, 1 4 of a full turn, 1 3 of a full turn, ]
Answer index: 0
Subject: Math
Skill: scalene-isosceles-and-equilateral-triangles
Description: Is this triangle scalene?
Picture: [Uncaptioned image]
Choices: [yes, no, ]
Answer index: 1
Subject: Math
Skill: select-figures-with-a-given-area